# Causal Explanation of Concept Drift –

### v1

@@ 1 Introduction
In the following, we consider a data space ${\mathcal{X}}$ composed of multiple features. We refer to the index set of all features as $\mathcal{F},|\mathcal{F}|<\infty$. Every feature $f\in\mathcal{F}$ takes values in the real numbers ${\mathbb{R}}$, i.e., ${\mathcal{X}}={\mathbb{R}}^{\mathcal{F}}$. For a subset $F\subset\mathcal{F}$ we write ${\mathcal{X}}_{F}$ for the subspace based on the features in $F$, and for a data point $X$, we write $X_{F}$ for the projection onto ${\mathcal{X}}_{F}$. In addition, if $P$ is a probability measure on ${\mathcal{X}}$ we also write $P(X_{F}=x)$ or $P_{|{\mathcal{X}}_{F}}$ for the marginalization of $P$ onto ${\mathcal{X}}_{F}$.
@@ 4 Computational Causality
such that $x\mapsto E_{F}(x,A)$ is measurable and $A\mapsto E_{F}(x,A)$ is a probability measure for all $x$ and $A$, respectively. We call $E_{F}(x)$ the $F$-manipulation of $E$ to $x$ and refer to $E_{\emptyset}(*)$ as the not-manipulated experiment. (Here we set ${}{\mathcal{X}}_{\emptyset}=\{*\}$)
$\>T\>$
@@
Algorithmically speaking, the objective of computational causality is to recover an underlying causal model from purely observational data. Concretely, given a dataset generated under the non-intervention experiment $E_{\emptyset}(*)$, we want to infer the complete causal structure over all possible interventions in $E$. In practical terms, this approach aims to bypass the need to conduct every intervention experimentally. Achieving this reconstruction requires imposing assumptions that guarantee 1) the existence of a valid causal model and 2) computational tractability of the inference procedure. Typical assumptions are the Markov property, full information, or faithfulness that constrain the distribution of the observed variables and the considered setup, making it more similar to setups covered by those produced by Bayesian networks. A detailed discussion of these assumptions is beyond the scope of this paper. For our experimental evaluation, we focus exclusively on the classical PC algorithm [21].
@@ 5.1 A Causal Model over Time
Following the ideas of [8, 11], that is, considering time as a feature with drift the dependency of time and data, allows us to easily model drift in the context of an experiment with intervention by simply extending the dataspace from ${\mathcal{X}}$ to ${\mathcal{X}}\times{\mathcal{T}}$, with the extended feature set ¯=$\overline{\mathcal{F}}=\mathcal{F}\sqcup\{{f_{\mathcal{T}}}\}$, where ${f_{\mathcal{T}}}\not\in\mathcal{F}$ represents the specific time feature and we also write $X_{{f_{\mathcal{T}}}}=T$ for consistency.
@@
[…]
@@ 5.3 A Refined Approach
In the setup of Theorem˜2, the smallest set of features that needs to be altered to ensure that we can obtain every time window distribution is exactly given by the set of all children of ${f_{\mathcal{T}}}$. In other words: for every window $W\subset{\mathcal{T}},\,P_{T}(W)>0$ there is a graphical model $(G_{W},P_{W,f})$ on ${\mathcal{X}}$ such that $P_{G}(X\in A\mid\textnormal{do}(T\in W))=P_{G_{W}}(A)$ and $P_{W,f}=P_{f}$ for all ${f_{\mathcal{T}}}\not\in\textnormal{pa}_{G}(f)$. Conversely, for every $f\in\mathcal{F}$ that has ${f_{\mathcal{T}}}$ as a parent, there exists a window $W$ such that $P_{W,f}\neq P_{f}$.
@@
2:  $G\leftarrow\textsc{DetermineDAG}(S)$triangleright Run causal discovery algorithm, e.g., PC
@@ Appendix A Proofs
Proof of Theorem˜2.  1. Follows directly from ${f_{\mathcal{T}}}$ being causeless.
2. In a faithful DAG a feature is a connected component in the graph if and only if the feature is independent of the rest. As $\textnormal{pa}({f_{\mathcal{T}}})=\emptyset$ we have $X\perp\!\!\!\perp T$ if and only if there is some $f\in\mathcal{F}$ such that ${f_{\mathcal{T}}}\in\textnormal{pa}(f)$.
@@
Choose any topological order $\sigma$ of $G$. Create $G^{\prime}$ by (i) adding $g\!\to\!f$ for all distinct $g,f\in C$ with $\sigma(g)<\sigma(f)$ and (ii) deleting the time node ${f_{\mathcal{T}}}$ and all its incident edges.
@@
$q_{f}\!\bigl{(}x_{f}\mid x_{\textnormal{pa}_{G^{\prime}}(f)}\bigr{)}:=\begin{cases}P_{G}\!\bigl{(}x_{f}\mid x_{\textnormal{pa}_{G^{\prime}}(f)},\,T\in W\bigr{)},&f\in C,\\[6.0pt] P_{G}\!\bigl{(}x_{f}\mid x_{\textnormal{pa}_{G^{\prime}}(f)}\bigr{)},&f\in O,\end{cases}$
@@
Since each kernel $q_{f}$ is indexed by $\textnormal{pa}_{G^{\prime}}(f)$ and their product equals $P_{G}(x\mid T\in W)$, the pair $(G^{\prime},Q)$ is a graphical model of the conditional distribution $P_{G}(X\in\,\cdot\mid T\in W)$. $\square$

### v2

@@ 1 Introduction
In the following, we consider a data space $\mathcal{X}$} composed of multiple features. We refer to the index set of all features as $\mathcal{F},|\mathcal{F}|<\infty$. Every feature $f\in\mathcal{F}$ takes values in the real numbers $\mathbb{R}$}, i.e., $\mathcal{X}$$\mathbb{R}$$^{\mathcal{F}}$. For a subset $F\subset\mathcal{F}$ we write ${\mathcal{X}}_{F}$ for the subspace based on the features in $F$, and for a data point $X$, we write $X_{F}$ for the projection onto ${\mathcal{X}}_{F}$. In addition, if $P$ is a probability measure on $\mathcal{X}$} we also write $P(X_{F}=x)$ or $P_{|{\mathcal{X}}_{F}}$ for the marginalization of $P$ onto ${\mathcal{X}}_{F}$.
@@ 4 Computational Causality
such that $x\mapsto E_{F}(x,A)$ is measurable and $A\mapsto E_{F}(x,A)$ is a probability measure for all $x$ and $A$, respectively. We call $E_{F}(x)$ the $F$-manipulation of $E$ to $x$ and refer to $E_{\emptyset}(*)$ as the not-manipulated experiment. (Here we set $\mathcal{X}$$_{\emptyset}$$\{*\}$)
T\>T\>
@@
Algorithmically speaking, the objective of computational causality is to recover an underlying causal model from purely observational data. Concretely, given a dataset generated under the non‐intervention experiment $E_{\emptyset}(*)$, we want to infer the complete causal structure over all possible interventions in $E$. In practical terms, this approach aims to bypass the need to conduct every intervention experimentally. Achieving this reconstruction requires imposing assumptions that guarantee 1) the existence of a valid causal model and 2) computational tractability of the inference procedure. Typical assumptions are the Markov property, full information, or faithfulness that constrain the distribution of the observed variables and the considered setup, making it more similar to setups covered by those produced by Bayesian networks. A detailed discussion of these assumptions is beyond the scope of this paper. For our experimental evaluation, we focus exclusively on the classical PC algorithm [21].
@@ 5.1 A Causal Model over Time
Following the ideas of [8, 11], that is, considering time as a feature with drift the dependency of time and data, allows us to easily model drift in the context of an experiment with intervention by simply extending the dataspace from $\mathcal{X}$} to ${\mathcal{X}}\times{\mathcal{T}}$, with the extended feature set $\overline{\mathcal{F}}=\mathcal{F}\sqcup\{{f_{\mathcal{T}}}\}$, where $f_{\mathcal{T}}$$\not\in\mathcal{F}$ represents the specific time feature and we also write $X_{{f_{\mathcal{T}}}}=T$ for consistency.
@@
${\mathcal{D}}_{t}\text{ has drift }\Longleftrightarrow E_{{f_{\mathcal{T}}}}\text{ is not constant ($P_{T}$-a.s.)}$
@@ 5.3 A Refined Approach
In the setup of Theorem˜2, the smallest set of features that needs to be altered to ensure that we can obtain every time window distribution is exactly given by the set of all children of $f_{\mathcal{T}}$}. In other words: for every window $W\subset{\mathcal{T}},\,P_{T}(W)>0$ there is a graphical model $(G_{W},P_{W,f})$ on $\mathcal{X}$} such that $P_{G}(X\in A\mid\textnormal{do}(T\in W))=P_{G_{W}}(A)$ and $P_{W,f}=P_{f}$ for all ${f_{\mathcal{T}}}\not\in\textnormal{pa}_{G}(f)$. Conversely, for every $f\in\mathcal{F}$ that has $f_{\mathcal{T}}$} as a parent, there exist two time windows $W,W^{\prime}$ such that $P_{W,f}\neq P_{W^{\prime},f}$.
@@
2:  $G\leftarrow\textsc{DetermineDAG}(S)$ $\triangleright$ Run causal discovery algorithm, e.g., PC
@@ Appendix A Proofs
Proof of Theorem˜2.  1. Follows directly from $f_{\mathcal{T}}$} being causeless. For any time window $W$ with positive measure and $A\in\Sigma_{\mathcal{F}}$, we have
$\int_{W}P(X\in S\mid T=x)\textnormal{d}P_{T}=P(X\in A\mid T\in W)$
and therefore $f_{\mathcal{T}}$} can be chosen as the first node in the graph node ordering of $G$, $f_{\mathcal{T}}$} has therefore no parents.
2. In a faithful DAG a feature is a connected component in the graph if and only if the feature is independent of the rest[21]. As $\textnormal{pa}({f_{\mathcal{T}}})=\emptyset$ we have $X\perp\!\!\!\perp T$ if and only if there is some $f\in\mathcal{F}$ such that ${f_{\mathcal{T}}}\in\textnormal{pa}(f)$.
@@
Choose any topological order $\sigma$ of $G$. Create $G^{\prime}$ by (i) adding $g\!\to\!f$ for all distinct $g,f\in C$ with $\sigma(g)<\sigma(f)$ and $f\in C$,
$P_{G}\!\bigl(x_{f}\mid x_{\textnormal{pa}_{G^{\prime}}(f)}\bigr)$, (ii) deleting the time node $f_{\mathcal{T}}$} and all its incident edges.
@@
$q_{f}\!\bigl(x_{f}\mid x_{\textnormal{pa}_{G^{\prime}}(f)}\bigr):=P_{G}\!\bigl(x_{f}\mid x_{\textnormal{pa}_{G^{\prime}}(f)},\,T\in W\bigr)$  $\qquad\textnormal{ for }f\in O,$
@@
Since each kernel $q_{f}$ is indexed by $\textnormal{pa}_{G^{\prime}}(f)$ and their product equals $P_{G}(x\mid T\in W)$, the pair $(G^{\prime},Q)$ is a graphical model of the conditional distribution $P_{G}(X\in\,\cdot\mid T\in W)$.
Regarding the minimality we fix $f\in\textnormal{ch}({f_{\mathcal{T}}})$. Let $U\subseteq\mathcal{F}\setminus\{f\}$ be the set of parents of $f$ in $G_{W}$, allowing any, but fixed, parents. Since $f$ is connected to $f_{\mathcal{T}}$} in $G$, a faithful graph, we have $X_{f}{\not\perp\!\!\!\perp}T\mid X_{U}$. We want to show, that there are $W,W^{\prime}\in\Sigma_{f_{\mathcal{T}}},A_{f}\in\Sigma_{f}$ and $A_{U}\in\Sigma_{U}$, such that
$0\neq\int_{A_{U}}(P_{W,f}(A_{f}|a_{U})-P_{W^{\prime},f}(A_{f}\mid a_{U}))\textnormal{d}P_{X_{U}}(a_{U})$
holds. Hence $X_{f}{\not\perp\!\!\!\perp}T\mid X_{U}$ there exist $A,W$, such that
$V(a_{U})$ $:=P(X_{f}\in A_{f}\mid X_{U}=a_{u})$
$U(a_{U})$ $:=P(T\in W\mid X_{U}=a_{U})$
$W(a_{U})$ $:=P(X_{f}\in A_{f},T\in W\mid X_{U}=a_{U})$
such that the delta $D(a_{u}):=W(a_{u})-V(a_{u})U(a_{u})$ is not 0 $P_{U}$-a.s., i.e. $0\not\equiv D$. W.l.o.g., we assume that $S^{+}:=\{a_{U}\in{\mathcal{X}}_{U}\mid D(a_{U})>0\}$ has a positive measure. Note, that on $S^{+}$ in addition $0<V(a_{U})<1$ holds almost surely.
By definition, we have
$P_{W,f}(X_{f}\in A_{f},X_{U}=a_{U})$ $=\frac{P(X_{f}\in A_{f}\mid T\in W,X_{U}=a_{U})}{P(X_{f}\in A_{f}\mid X_{U}=a_{u})}=U(a_{u})+\frac{D(a_{u})}{V(a_{u})}$
$P_{W^{C},f}(A_{f},a_{U})$ $=\frac{P(X_{f}\in A_{f}\mid T\in W^{C},X_{U}=a_{U})}{P(X_{f}\in A_{f})}=U(a_{U})-\frac{D(a_{U})}{1-V(a_{U})}$
(which is well defined on $S^{+}$) and therefore, the difference of the kernels
$P_{W,f}(A_{f},a_{U})-P_{W^{C},f}(A_{f},a_{U})=\frac{D(a_{U})}{V(a_{U})(1-V(a_{U}))}$
is positive on $S^{+}$, the kernels are not a.s. equal.
$\square$